# FlowDubber: Movie Dubbing with LLM-based Semantic-aware Learning and Flow Matching based Voice Enhancing

### v1

@@ 4.4. Comparison with SOTA Dubbing Methods
To prove the effectiveness of the proposed model, we compare the proposed method with the most advanced dubbing models. All experimental results use the official code or providing checkpoint.
Results on the Chem Dataset. As shown in Table 1, our method achieves the best performance on almost all metrics on the Chem benchmark, whether in setting or setting 2. First, our method achieves the best LSE-C and LSE-D, with absolute improvements of 5.63% and 5.65% than the SOTA dubbing method ProDubber (Zhang et al., 2025a), demonstrating the effectiveness of our methods in lip-sync by LLM-based semantic-aware learning and dual contrastive aligning. Furthermore, in the speaker similarity (see SIM-O), our method improves 13.72% on dub setting1 and 11.15% on dub setting2 than Speaker2Dubber (Zhang et al., 2025a). Please note that setting 2 is more challenging than setting 1. Besides, the dubbing synthesis quality of our method is the highest among all dubbing methods, with a UTMOS score of 3.91. In summary, FlowDubber  is a comprehensive dubbing model that makes up for the shortcomings of previous methods in audio-visual synchronization, speaker similarity, and dubbing synthesis quality, and achieves a WER comparable to SOTA.
Results on the GRID Dataset. As shown in Table 3, a similar trend is found in the multi-speaker benchmark. We still achieve SOTA performance in audio-visual synchronization, dubbing synthesis quality, and discrepancy from ground truth in both dubbing settings while maintaining similarity with advanced speaker identity. Specifically, our method can achieve similar WER as ProDubber (Zhang et al., 2025a) while maintaining high audio-visual synchronization: ours LSE-C is absolute improved by 2.03% and LSE-D is absolute improved by 2.87% compared with Produbber, which is extremely important for moving towards automated and high-quality dubbing. Besides, the UTMOS of our method is improved by 12% over Speaker2Dubber (Zhang et al., 2024) on setting 2, which shows that the speech quality synthesized by our dubbing method is the best, even better than the two-stage pre-training manner, which benefits from the proposed voice enhanced flow matching and LLM-based semantic-aware learning.
Results on the Speaker Zero-shot Test. In addition to dubbing benchmarks, we also conduct the zero-shot test to evaluate the generalization performance of models. This setting uses the audio of unseen characters (from another dataset) as reference audio. Here, we use the audio from the Chem dataset as reference audio to measure the GRID dataset. Since there is no target audio at this setting, we only compare LSE-C, LSE-D, WER, and UTMOS for objective evaluation. As shown in Table 2, our proposed method surpasses the current state-of-the-art models and achieves the best performance across all metrics. Specifically, our proposed method achieves the best pronunciation accuracy 13.96% and the best acoustic quality 3.98 than SOTA dubbing method Produbber (Zhang et al., 2025a), even facing the out-of-domain reference audio. Besides, we still achieve the best lip-sync (see LSE-C and LSE-D) in zero-shot setting, which proves the the superiority of the generalization performance of our proposed method.
@@
4.5. Analysis of Voice-Enhanced Flow Matching
To evaluate the effectiveness of LLM-based Acoustics Flow Matching Guidance in improving the clarity of generated speech, we assess the DNSMOS (Reddy et al., 2021), SNR Score and UTMOS metrics across different guidance scale. Table 4 shows the results. As the guidance scale increases, DNSMOS, SNR Score and UTMOS all show improvement, indicating that LLM-based Acoustics Flow Matching Guidance effectively reduces noise and enhances speech clarity, naturalness, and overall quality. Higher scales correlate with better noise suppression, intelligibility, and a more refined listening experience. Besides, DNSMOS increases faster than UTMOS, indicating that LLM-based Acoustics Flow Matching Guidance primarily enhances clarity, as DNSMOS is more closely associated with speech clarity than UTMOS.
Figure 3. The visualization of the mel-spectrograms of ground truth (GT) and synthesized audios obtained by different models. In (a), green arrows point to the video frames that no speak, and green bounding boxes are used to highlight the pauses in the speech. In (b), pink arrows point to the enhanced details of the mel-spectrogram as flow matching guidance scale $\alpha$ increases.
@@ 4.6. Ablation Studies
To further investigate the specific effects of main module (LLM-SL, DCA, FVE, Style in FVE) in our proposed method, we conduct ablation studies on the Dub 1.0 setting of the Chem benchmark. The ablation results are presented in Table 5. It shows that all modules contribute significantly to the overall performance, and each module has a different focus. Specifically, when FVE is removed (line 1), UTMOS drops the most, which shows the importance of the proposed voice enhanced flow matching to gradually remove noise and generate high-quality mel-spectrogram. When LLM-SL (line 2)is removed, both WER and UTMOS decrease, with WER being more obvious. This shows that LLM-based semantic-aware learning can provide rich semantic information on phoneme level, which is necessary for clear pronunciation. When removing DCA and using the duration predictor (line 3) to provide alignment, we observe a significant degradation in LSE-C and LSE-D. Although the impact on sound quality is very small (see UTMOS), it is unacceptable for video dubbing. Last, removing Style in FVE has a greater impact on speaker similarity (see SIM-O).
Table 6. Compared with different audio generators. All results build on acoustics flow matching guidance of scale $\alpha$=0.8.
@@
4.7. Compare with Different Audio Generators
Please note that when comparing with the dubbing baseline (Table 1324, and 5), we adopt HiFi-GAN (Kong et al., 2020) as audio generator to convert the mel-spectrogram to waveforms for ensure fairness. Although HiFi-GAN is one of the most popular vocoders, it is not the most advanced. Thus, we explore the upper-bound quality of the generated audio in this section by using different audio generators. Specifically, we select more powerful audio generators: BigVGAN (Lee et al., 2023b), 16K Hz Descript Audio Codec (DAC) (Kumar et al., 2023), and 24K Hz Codec Vocoder (CV) (Du et al., 2024b), respectively. Note that since the DAC and CV do not accept features in the form of mel-spectrogram, we first convert the mel-spectrogram into a waveform by basic HiFi-GAN, then use their codec manner to discretize it, and finally reconstruct the target audio. As shown in Table 6, the results show that 24K CV achieves the best speech quality (see UTMOS), while BigVGAN achieves better alignment and timbre restoration with a slight advantage. Most importantly, we find that all audio generators are better than SOTA dubbing baseline (e.g., Produbber (Zhang et al., 2025a)) or powerful TTS methods (see Table 7) in audio-visual synchronization (see LSE C/D), because the aligning information has been preserved in advance. This is also the advantage of our method, which can be extended by stronger audio generators in the future.
4.8. Compare with LLM-based TTS method
As shown in Table 7, we compare with the recent LLM-based TTS methods. Our method achieves the best performance in LSE-C and LSE-D to maintain synchronization, while ensuring high speech quality. Specifically, our dubbing scheme can approach or even exceed some large-scale TTS methods in UTMOS. For example, our UTMOS is 3.59% higher than FireRedTTS. In contrast, LLM-based TTS methods cannot adapt to dubbing scenes due to the lower LSE-D and LSE-C, proving the bad audio-visual alignment with lip motion. Although Spark-TTS has excellent speech quality, its speaker similarity is poor. In terms of speaker similarity, our method can improve 31.32% on SIM-O than Spark-TTS.
4.9. Qualitative Analysis
We visualize the mel-spectrograms of ground truth and dubbing generated by different models for comparison in Figure 3. The green bounding boxes highlight the pauses in the speech and blue bounding boxes exhibit significant differences in acoustic details. We have also enlarged the details to make it easier for readers to compare. As shown in Figure 3 (a), our method demonstrates high-quality audio-visual alignment and acoustic quality relative to state-of-the-art dubbing baseline. In the corresponding silent video frames (see green arrows), our method can generate the same sound pauses as GT, which illustrates the effectiveness of dual contrastive aligning. As shown in Figure 3 (b), we visualize the mel-spectrogram generation effect of Acoustics Flow Matching Guidance at different scales. As the scale increases, the originally blurry and artifact-filled spectrum gradually becomes clearer. The qualitative analysis shows that our model can generate high-quality audio-visual alignment, high-fidelity acoustic quality and speech.
@@ 5. Conclusion
In this paper, we propose an LLM-based dubbing architecture, which incorporates speech large language model for semantic-aware learning and voice enhanced flow matching for high-quality acoustic modeling. By LLM-based semantic-aware learning, the model absorbs the phoneme level semantic knowledge with in-contextual information, while maintaining the lip-sync with silent video by dual contrastive aligning. Besides, the proposed Flow-based Voice Enhancing ensures the acoustic clarity and speaker identity by LLM-based Acoustics Flow Matching Guidance and style flow matching prediction. Our proposed model sets new state-of-the-art on both Chem and GRID benchmarks.
Appendix A Theoretic Details of Flow Matching
Given the mel-spectrograms data space with data point $M$, where $M\sim q(M)$ and $q(M)$ is an unknown data distribution of mel-spectrograms, a possible approach to sample $M$ from $q(M)$ is from the probability density path defined as $p_{t}(x)$ where $t\in[0,1]$, $p_{0}(x)=\mathcal{N}(x;\boldsymbol{0},\boldsymbol{I})$ and $p_{1}(x)\approx q(x)$. To estimate the probability density path, Continuous Normalizing Flow (Chen et al., 2018) defines a vector field $v_{t}$, which gives a flow $\phi_{t}(x)$ through an Ordinary Differential Equation (ODE):
(13) $\frac{d}{dt}\phi_{t}(x)=v_{t}(\phi_{t}(x));\quad\phi_{0}(x)=x.$
Chen et al. (2018) shows that predicting the vector flow $v_{t}$ through a neural network $\theta$ can be used to transform a simple Gaussian distribution to a more complicated one, such as $q(x)$, which generates the probability density path. Flow matching is designed to predict such a probability density path. Give a target probability density path $p_{t}$ with its corresponding known vector field $u_{t}(x)$, the training objective of flow matching is:
(14) $\mathcal{L}_{FM}(\theta)=\mathbb{E}_{t,p_{t}(x)}||v_{t}(x,\theta)-u_{t}(x)||^{2}$
However, the total probability density path $p_{t}$ is unknown and we can only use some samples from $q(x)$ to estimate the probability density path, which is Conditional Flow Matching (CFM). The training objective of CFM is:
(15) $\mathcal{L}_{CFM}(\theta)=\mathbb{E}_{t,M\sim q(M),p_{t}(x|M)}||v_{t}(x,\theta)-u_{t}(x|M)||^{2},$
We can efficiently estimate the probability density path $p_{t}(x)$ by sampling from $q(M),p_{t}(x|M)$ and calculate $u_{t}(x|M)$. Here we use optimal-transport conditional flow matching (OT-CFM) to train our model, which is a simple version of CFM with simple flow $\phi_{t}(x)=(1-(1-\sigma_{\min})t)x_{0}+tM$, which satisfies $x_{0}\sim\mathcal{N}(x;\boldsymbol{0},\boldsymbol{I})$ and $\phi_{1}(x)\sim q(x)$. Its gradient vector field is $u_{t}(\phi_{t}(x)|M)=M-(1-\sigma_{min})x_{0}$, enabling fast training and inference for its linear and time-invariant properties. Given conditional mean $\mu$, the training objective of OT-CFM can be formulated as:
(16) $\mathcal{L_{\theta}}=\mathbb{E}_{t,q(M),p_{t}(x|\mu,M)}||v_{t}(\phi_{t}(x)|\mu,\theta)-u_{t}(\phi_{t}(x)|M)||^{2},$
where $v_{t}(\phi_{t}(x)|\mu,\theta)$ is the predicted gradient vector field of $\phi_{t}(x)$ according to the acoustics prior information $\mu$. Then, we can solve the ODE $d\phi_{t}(x)=v_{t}(\phi_{t}(x)|\mu,\theta)dt$ from $t=0$ to $t=1$ to generate the target mel-spectrogram $\hat{M}$ from noise $x_{0}$.
Table 8. Subjective evaluation on GRID benchmark.
Dataset GRID
Methods MOS-N $\uparrow$ MOS-S $\uparrow$ CMOS $\uparrow$
GT 4.69$\pm$0.07 - +0.10
V2C-Net (Chen et al., 2022) 3.62$\pm$0.06 3.67$\pm$0.11 -0.35
HPMDubbing (Cong et al., 2023) 3.77$\pm$0.20 3.74$\pm$0.13 -0.26
StyleDubber (Cong et al., 2024b) 4.02$\pm$0.11 4.06$\pm$0.05 -0.19
Speaker2Dubber (Zhang et al., 2024) 4.10$\pm$0.09 4.05$\pm$0.11 -0.18
Produbber (Zhang et al., 2025a) 4.12$\pm$0.07 4.07$\pm$0.10 -0.13
Our method 4.15$\pm$0.06 4.10$\pm$0.07 0.00
Appendix B LSE or MCD-DTW-SL for measuring lip-sync?
MCD-DTW-SL cannot truly measure audiovisual synchronization because the coefficients of MCD-DTW-SL are based on the global time rather than the fact to reflect the alignment related to lip movement. Below is the formula of MCD-DTW-SL:
[…]
(18) $n=\frac{T_{mel}}{T_{v}}=\frac{sr/hs}{FPS}\in\mathbb{N}^{+},$
where $FPS$ denotes the Frames per Second of the video, $sr$ denotes the sampling rate of the audio, and $hs$ denotes hop size when transforming the raw waveform into mel-spectrograms. In this case, the audio length is known in advance, MCD-DTW-SL is meaningless for determining alignment. Thus, we encourage evaluating the audio-visual synchronization in movie dubbing by using metrics LSE-D and LSE-C, which are widely adopted for quantitative evaluation of lip-syncing performance in the wild (Yaman et al., 2024; Ki and Min, 2023; Yang et al., 2024a; Tan et al., 2024; Park et al., 2024). The LSE-D measures the distance between the audio and visual representations with lower scores suggesting better audio-visual sync. LSE-C is the confidence score, and the higher value implies a stronger correlation between video and speech (Lei et al., 2024; Liu et al., 2024; Sahipjohn et al., 2024).
Appendix C Subjective Evaluation on GRID Benchmark
Since we only provide objective evaluation in the main paper, we provide the subjective evaluation by human, following previous dubbing works (Zhang et al., 2024, 2025a). Please note that the UTMOS and DNSMOS in the paper are given by the model prediction to ensure fairness, rather than human subjective evaluation.
MOS-N & MOS-S. MOS-Naturalness (MOS-N) and MOS-Similarity (MOS-S) are mean option scores reported with a 95% confidence interval based on ratings from 20 native English speakers using a scale from 1 to 5. Each participant is required to listen to 30 randomly selected generated dubbing and rate the dubbing according to the speech naturalness and voice similarity following (Chen et al., 2022).
CMOS. Comparative mean option score (CMOS) asks participants to compare the dubbing generated by two models using same input and rate them on a scale from -5 to 5 based on criterion of matched degree between generated dubbing with video (Zhang et al., 2025a).
Under the same experimental setting, we found that the proposed FlowDubber is also subjectively superior to the previous methods, especially in speech quality (MOS-N), indicating the best overall dubbing quality.

### v2

@@ 4.4. Comparison with SOTA Dubbing Methods
Results on the Chem Dataset. As shown in Table 1, our method achieves the best performance on almost all metrics on the Chem benchmark, whether in setting 1 or setting 2. First, our method achieves the best LSE-C and LSE-D, with absolute improvements of 5.63% and 5.65% than the TTS-based dubbing methods with duration predictor (like StyleDubber (Cong et al., 2024b), Speaker2Dubber (Zhang et al., 2024b), Produbber (Zhang et al., 2025a)), demonstrating the effectiveness of our methods in lip-sync by LLM-based semantic-aware learning and dual contrastive aligning. Besides, the dubbing synthesis quality of our method is the highest among all dubbing methods, with a UTMOS score of 3.91. In summary, FlowDubber  is a comprehensive dubbing model that makes up for the shortcomings of previous methods in audio-visual synchronization, speaker similarity, and dubbing synthesis quality, and achieves a WER comparable to SOTA.
Results on the GRID Dataset. As shown in Table 3, a similar trend is found in the multi-speaker benchmark. We still achieve SOTA performance in audio-visual synchronization, dubbing synthesis quality, and discrepancy from ground truth in both dubbing settings while maintaining similarity with advanced speaker identity. Specifically, our method can achieve similar WER as ProDubber (Zhang et al., 2025a) while maintaining higher LSE-C/D than previous TTS-based dubbing methods (like StyleDubber, Speaker2Dubber, Produbber), which adopt a Duration Predictor (DP) to produce rough duration, leading to poor audio-visual alignment. Finally, the UTMOS of our method is improved by 12% over Speaker2Dubber (Zhang et al., 2024b) on setting 2, which shows that the speech quality synthesized by our method is the best, even better than the two-stage pre-training manner.
Results on the Speaker Zero-shot Test. In addition to dubbing benchmarks, we also conduct the zero-shot test to evaluate the generalization performance of models. This setting uses the audio of unseen characters (from another dataset) as reference audio. Here, we use the audio from the Chem dataset as reference audio to measure the GRID dataset. As shown in Table 2, our proposed method surpasses the current state-of-the-art models and achieves the best performance across all metrics. Besides, we still achieve the best lip-sync (see LSE-C and LSE-D) in zero-shot setting.
@@
4.5. Analysis of Flow-based Voice Enhancing
As shown in 4, we use DNSMOS, SNR, and UTMOS as main metrics. As the guidance scale increases, DNSMOS, SNR, and UTMOS all show improvement, indicating that LLM-based Acoustics Flow Matching Guidance effectively reduces noise and enhances speech clarity and overall intelligibility. Besides, we find that DNSMOS increases faster than UTMOS, indicating that the proposed method primarily enhances clarity.
Figure 3. The visualization of the mel-spectrograms of ground truth (GT) and synthesized audios obtained by different models. In (a), green arrows point to the video frames that do not speak, and green bounding boxes are used to highlight the pauses in speech. In (b), pink arrows point to the enhanced details of the mel-spectrogram as flow matching guidance scale $\alpha$ increases.
@@ 4.6. Ablation Studies
To further investigate the specific effects of main module in our method, we conduct ablation studies on the Dub 1.0 setting of the Chem benchmark. The ablation results are presented in Table 5. It shows that all modules contribute significantly to the overall performance, and each module has a different focus. When LLM-SL is removed, both WER and UTMOS decrease, with WER being more obvious. This shows that LLM-based semantic-aware learning can provide rich semantic information on phoneme level, which is necessary for clear pronunciation. When removing DCA and using the duration predictor to provide alignment, we observe a significant degradation in LSE-C and LSE-D. Last, removing Style in FVE has a greater impact on speaker similarity (see SIM-O).
4.7. Compare with Different Audio Generators
Please note that when comparing with the dubbing baseline (Table 1-5), we adopt HiFi-GAN (Kong et al., 2020) as audio generator to convert the mel-spectrogram to waveforms to ensure fairness. To explore the upper-bound quality of the generated audio by using different audio generators, we select more powerful audio generators: BigVGAN (Lee et al., 2023b), 16K Hz Descript Audio Codec (DAC) (Kumar et al., 2023), and 24K Hz Codec Vocoder (CV) (Du et al., 2024b), respectively. To ensure the integrity of the original design (without removing the FVE module), we do not consider directly decoding waveforms from tokens. Therefore, for DAC and CV, we first generate the original waveform and then perform reconstruction. As shown in Table 6, the results show that 24K CV achieves the best speech quality (see UTMOS), while BigVGAN achieves better alignment and timbre restoration with a slight advantage. Most importantly, we find that all audio generators are better than SOTA dubbing baseline (e.g., Produbber (Zhang et al., 2025a)) or powerful TTS methods (see Table 7) in audio-visual synchronization (see LSE C/D), because the aligning information has been preserved in advance. This is also the advantage of our design, which can be extended by stronger audio generators in the future.
Table 6. Compared with different audio generators. The results under the $\alpha$=0.8 guidance scale of FVE.
@@ 4.7. Compare with Different Audio Generators
4.8. Compare with LLM-based TTS method
As shown in Table 7, we compare with the recent LLM-based TTS methods. Our method achieves the best performance in LSE-C and LSE-D to maintain synchronization, which is extremely important for moving towards automated lip-sync dubbing. Besides, our dubbing scheme can approach or even exceed part of large-scale TTS methods in UTMOS. For example, our UTMOS is 3.59% higher than FireRedTTS. In contrast, most part of LLM-based TTS methods cannot adapt to dubbing scenes due to the lower LSE-D and LSE-C, proving the bad audio-visual alignment with lip movement.
4.9. Qualitative Analysis
We visualize the mel-spectrograms of ground truth and dubbing generated by different models for comparison in Figure 3. The green bounding boxes highlight the pauses in the speech, and blue bounding boxes exhibit significant differences in acoustic details. We have also enlarged the details to make it easier for readers to compare. As shown in Figure 3(a), our method demonstrates high-quality audio-visual alignment and acoustic quality relative to state-of-the-art dubbing baseline. In the corresponding silent video frames (see green arrows), our method can generate the same sound pauses as GT, which illustrates the effectiveness of dual contrastive aligning. As shown in Figure 3(b), we visualize the mel-spectrogram generation effect of Acoustics Flow Matching Guidance at different scales. As the scale increases, the originally blurry and artifact-filled spectrum gradually becomes clearer. The qualitative analysis shows that our model can generate high-quality audio-visual alignment and high-fidelity acoustic quality.
@@ 5. Conclusion
In this paper, we propose an LLM-based dubbing architecture, which incorporates a large language model for semantic-aware learning and voice-enhanced flow matching for acoustic modeling. By LLM-based semantic-aware learning, the model absorbs the phoneme-level semantic knowledge with in-contextual information, while maintaining the lip-sync by dual contrastive aligning. Besides, the flow-based voice enhancing ensures the acoustic clarity and speaker identity. Our proposed model sets SOTA results on both Chem and GRID benchmarks. In the future, we will explore the wild datasets and provide lightweight solutions to perform fast inference.
Acknowledgments
This work was supported by the National Nature Science Foundation of China (62322211), the ”Pioneer” and ”Leading Goose” R&D Program of Zhejiang Province (2024C01023), Key Laboratory of Intelligent Processing Technology for Digital Music (Zhejiang Conservatory of Music), Ministry of Culture and Tourism (2023DMKLB004). Amin Beheshti, Anton van den Hengel, and Yuankai Qi are not supported by the aforementioned fundings.